# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v2

@@ 2 Datasets and Benchmarks
 (§4.3), fill=yellow!20 [ Context-Enrichment (§4.3.1), fill=red!15 [ EMERGE Zhu et al. (2024b), MiRAG Adjali et al. (2024), Wiki-LLaVA Caffagni et al. (2024), Video-RAG Luo et al. (2024b), Img2Loc Zhou et al. (2024e),  Xue et al. (2024) , leaf, text width=62.0em ] ] [ Adaptive and
Iterative
Retrieval (§4.3.2), fill=red!15 [ SKURG Yang et al. (2023), IRAMIG Liu et al. (2024b), OMG-QA Nan et al. (2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al. (2024a), OmniSearch Li et al. (2024b), mR2AG Zhang et al. (2024f), RAGAR Khaliq et al. (2024) , leaf, text width=62.0em ] ] ] [ Generation
Technique
s (§4.4), fill=yellow!20 [ In-Context
Learning (§4.4.1), fill=red!15 [ RMR Tan et al. (2024),  Sharifymoghaddam et al. (2024), RA-CM3 Yasunaga et al. (2023), RAG-Driver Yuan et al. (2024), MSIER Luo et al. (2024a), Raven Rao et al. (2024) , leaf, text width=62.0em ] ] [ Reasoning (§4.4.2), fill=red!15 [ RAGAR Khaliq et al. (2024), VisDoM Suri et al. (2024), SAM-RAG Zhai (2024), LDRE Yang et al. (2024) , leaf, text width=62.0em ] ] [ Instruction
Tuning (§4.4.3), fill=red!15 [ RA-BLIP Ding et al. (2024b), RAGPT Lang et al. (2025), mR2AG Zhang et al. (2024f), RagVL Chen et al. (2024d), Jang et al. (2024), MMed-RAG Xia et al. (2024a), MegaPairs Zhou et al. (2024a), Surf Sun et al. (2024a), Rule Xia et al. (2024b) , leaf, text width=62.0em ] ] [ Source
Attribution (§4.4.4), fill=red!15 [ MuRAR Zhu et al. (2025), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=62.0em ] ] ] [ Training
@@
 (§4.5), fill=yellow!20 [ Alignment (§4.5.1), fill=red!15 [ VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al. (2024),  Zhi Lim et al. (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al. (2024) , leaf, text width=62.0em ] ] [ Robustness (§4.5.3), fill=red!15 [  Buettner and Kovashka (2024), MORE Cui et al. (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al. (2024), RA-BLIP Ding et al. (2024b), RagVL Chen et al. (2024d), RA-CM3 Yasunaga et al. (2023) , leaf, text width=62.0em ] ] ] ] ]
@@
3 Evaluation
Evaluating multimodal RAG models is complex due to their varied input types and complex structure. The evaluation combines metrics from VLMs, generative AI, and retrieval systems to assess capabilities like text/image generation and information retrieval. Our review found about 60 different metrics used in the field. More details, including the formulas for the RAG evaluation metrics, can be found in Appendix (§C). In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
Retrieval Evaluation
Retrieval performance is measured through accuracy, recall, and precision metrics, with an F1 score combining recall and precision. Recall@K, which examines relevant items in top K results, is preferred over standard recall. Mean Reciprocal Rank (MRR) serves as another key metric for evaluation, which is utilized by Adjali et al. (2024); Nguyen et al. (2024).
Modality Evaluation
Modality-based evaluations primarily focus on text and image, assessing their alignment, text fluency, and image caption quality. For text evaluation, metrics include Exact Match (EM), BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005). MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings.
For image captioning, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al. (2015) measures caption quality using TF-IDF and cosine similarity Yasunaga et al. (2023); Zhao et al. (2024); Luo et al. (2024a); Yuan et al. (2024); Sharifymoghaddam et al. (2024); Hu et al. (2023); Rao et al. (2024); Xu et al. (2024a); Kim et al. (2024); Zhang et al. (2024c), while SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al. (2016) focuses on semantics. SPIDEr Liu et al. (2017), used in Zhang et al. (2024c), combines both metrics. For semantic alignment, BERTScore Zhang et al. (2020) compares BERT embeddings Sun et al. (2024b); Shohan et al. (2024), and evaluates fluency Chen et al. (2022a); Zhi Lim et al. (2024); Ma et al. (2024c). CLIP Score Hessel et al. (2021), used in Sharifymoghaddam et al. (2024); Zhang et al. (2024c), measures image-text similarity using CLIP Radford et al. (2021). For image quality, FID (Fréchet Inception Distance) Heusel et al. (2017) compares feature distributions Yasunaga et al. (2023); Zhao et al. (2024); Sharifymoghaddam et al. (2024); Zhang et al. (2024c), while KID (Kernel Inception Distance) Bińkowski et al. (2018) provides an unbiased alternative. Inception Score (IS) evaluates image diversity and quality through classification probabilities Zhi Lim et al. (2024). For audio evaluation, Zhang et al. (2024c) uses human annotators to assess sound quality (OVL) and text relevance (REL), while also employing Fréchet Audio Distance (FAD) Kilgour et al. (2019), an audio-specific variant of FID.
@@ 4.1 Retrieval Strategy
4.1.1 Efficient Search and Similarity Retrieval
Modern multimodal RAG systems encode diverse input modalities into a unified embedding space to enable direct cross-modal retrieval. Recent advancements in CLIP-based Radford et al. (2021) and BLIP-inspired Li et al. (2022a) approaches have driven the evolution of contrastive learning strategies through novel multimodal retrieval architectures and training methodologies Zhou et al. (2024c); Wei et al. (2024b); Zhang et al. (2024i). As these multi-encoder models project different modalities into a shared latent space, multimodal RAGs rely on efficient search strategies to retrieve relevant external knowledge.
Maximum inner product search (MIPS) variants are widely used for fast and direct similarity comparisons Tiwari et al. (2024); Wang et al. (2024c); Zhao et al. (2023b). Systems such as MuRAG Chen et al. (2022a) and RA-CM3 Yasunaga et al. (2023) employ approximate MIPS to efficiently retrieve the top-k candidates by maximizing the inner product between the query vector and a large collection of image–text embeddings. Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al. (2022), for high-speed retrieval. Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al. (2020), MAXSIM score Chan and Ng (2008); Cho et al. (2024), and approximate KNN methods Caffagni et al. (2024) have been adopted for efficient similarity computation.
Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy. These include adaptive quantization Zhang et al. (2023a); Li et al. (2024a), hybrid sparse-dense representations Nguyen et al. (2024); Zhang et al. (2024a), and learned index structures Zhai et al. (2023); Basnet et al. (2024). These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
4.1.2 Modality-Based Retrieval
Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality.
Text-Centric Retrieval
This remains foundational in multimodal RAG systems, with both traditional methods like BM25 Robertson and Zaragoza (2009), and dense retrievers such as MiniLM Wang et al. (2020a) and BGE-M3 Chen et al. (2024b) dominate text-based evidence retrieval Chen et al. (2022b); Suri et al. (2024); Nan et al. (2024). Novel approaches also address the need for fine-grained semantic matching and domain specificity: For instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al. (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al. (2024h) and CRAG Yan et al. (2024) enhance retrieval performance by ensuring precise citation of text spans.
Vision-Centric Retrieval
This focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al. (2023). Systems such as EchoSight Yan and Xie (2024) and ImgRet Shohan et al. (2024) retrieve visually similar content by using reference images as queries. In addition, composed Image Retrieval (CMI) models Feng et al. (2023); Zhao et al. (2024); Jang et al. (2024) enhance retrieval by integrating multiple image features into a unified query representation. Similarly, Pic2word Saito et al. (2023) maps visual content to textual descriptions, enabling zero-shot image retrieval.
Video-Centric Retrieval
These methods extend vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al. (2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al. (2024), optimizing multimodal transfer learning for video-text retrieval. Recent breakthroughs focus on long-context processing: Video-RAG Luo et al. (2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, T-Mass Wang et al. (2024b) models text as a stochastic embedding to enhance text-video retrieval, and VideoRAG Ren et al. (2025) -employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos. For temporal reasoning, CTCH Shen et al. (2024) uses contrastive transformer hashing to model long-term dependencies, while RTime Du et al. (2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality. Meanwhile, for complex video understanding, OmAgent Zhang et al. (2024e) adopts a divide-and-conquer framework, and DRVideo Ma et al. (2024d) addresses long-video understanding with a document-based retrieval approach.
@@ 4.1.2 Modality-Based Retrieval
Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages. These approaches integrate textual, visual, and structural elements like tables, font styles, and page layouts to enhance retrieval performance in complex documents. ColPali Faysse et al. (2024) pioneers end-to-end document image retrieval by embedding page patches with a vision-language backbone, bypassing OCR entirely. Models like ColQwen2 Wang et al. (2024d); Khattab and Zaharia (2020) and M3DocVQA Cho et al. (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning.
@@
Newer frameworks refine efficiency and layout understanding: ViTLP Mao et al. (2024) and DocLLM Wang et al. (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al. (2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computation costs. Finally, mPLUG-DocOwl 1.5 Hu et al. (2024a) and 2 Hu et al. (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies.
4.1.3 Re-ranking and Selection Strategies
Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also prioritizing retrieved candidates. Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms.
Optimized Example Selection
These techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al. (2024a); Yuan et al. (2023). For instance, Su et al. (2024a) enhance multimodal inputs using probabilistic control keywords to improve credibility, RULE Xia et al. (2024b) calibrates retrieved context selection via statistical methods like the Bonferroni correction to mitigate factuality risks, and clustering-based key-frame selection ensures diversity in video-based retrieval Dong et al. (2024b).
Relevance Score Evaluation
Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al. (2025b, a); Zhi Lim et al. (2024). Multimodal similarity measures, including structural similarity index measure (SSIM) Wang et al. (2020b), normalized cross-correlation (NCC), and BERTScore Zhang et al. (2020), aid in re-ranking retrieved documents. Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al. (2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al. (2024g); Yan and Xie (2024); Xu et al. (2024a). LDRE Yang et al. (2024) employs semantic ensemble methods to adaptively weigh multiple caption features, whereas RAGTrans Cheng et al. (2024) and OMG-QA Nan et al. (2024) incorporate traditional ranking functions like BM25 Robertson and Zaragoza (2009).
Filtering Mechasnism
This ensures high-quality retrieval by eliminating irrelevant data. Hard negative mining, as used in GME Zhang et al. (2024i) and MM-Embed Lin et al. (2024a), mitigates modality bias through modality-aware sampling and synthesized negatives. Similarly, consensus-based filtering approaches, seen in MuRAR Zhu et al. (2025) and ColPali Faysse et al. (2024), employs source attribution and multi-vector mapping to filter out low-similarity candidates. Moreover, dynamic modality filtering methods, such as RAFT Zhang et al. (2024h), Img2Loc Zhou et al. (2024e), and MAIN-RAG Chang et al. (2024), train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.

### v3

@@ 2 Datasets, Evaluation, and Applications
 (§3.3), fill=yellow!20 [ Context-Enrichment (§3.3), fill=red!15 [ EMERGE Zhu et al. (2024b), MiRAG Adjali et al. (2024), Wiki-LLaVA Caffagni et al. (2024), Video-RAG Luo et al. (2024b), Img2Loc Zhou et al. (2024e),  Xue et al. (2024b) , leaf, text width=62.0em ] ] [ Adaptive and Iterative
Retrieval (§3.3), fill=red!15 [ SKURG Yang et al. (2023), IRAMIG Liu et al. (2024b), OMG-QA Nan et al. (2024b), SAM-RAG Zhai (2024), MMed-RAG Xia et al. (2024a), OmniSearch Li et al. (2024d), mR2AG Zhang et al. (2024f), RAGAR Khaliq et al. (2024), UniversalRAG Yeo et al. (2025), OMGM Yang et al. (2025) , leaf, text width=62.0em ] ] ] [ Generation
Techniques
 (§3.4), fill=yellow!20 [ In-Context Learning  (§3.4), fill=red!15 [ RMR Tan et al. (2024),  Sharifymoghaddam et al. (2024), RA-CM3 Yasunaga et al. (2023), RAG-Driver Yuan et al. (2024), MSIER Luo et al. (2024a), Raven Rao et al. (2024) , leaf, text width=62.0em ] ] [ Reasoning (§3.4), fill=red!15 [ RAGAR Khaliq et al. (2024), VisDoM Suri et al. (2024), SAM-RAG Zhai (2024), LDRE Yang et al. (2024b) , leaf, text width=62.0em ] ] [ Instruction Tuning  (§3.4), fill=red!15 [ RA-BLIP Ding et al. (2024b), RAGPT Lang et al. (2025), mR2AG Zhang et al. (2024f), RagVL Chen et al. (2024e), Jang et al. (2024), MMed-RAG Xia et al. (2024a), MegaPairs Zhou et al. (2024a), Surf Sun et al. (2024a), Rule Xia et al. (2024b) , leaf, text width=62.0em ] ] [ Source Attribution  (§3.4), fill=red!15 [ MuRAR Zhu et al. (2025), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024b) , leaf, text width=62.0em ] ] [ Agentic Generation and Interaction  (§3.4), fill=red!15 [ AppAgent v2  Li et al. (2024c), USER-LLM R1 Rahimi et al. (2025), MMAD Jiang et al. (2025),  Yi et al. (2025), CollEX Schneider et al. (2025), HM-RAG Liu et al. (2025a), CogPlanner Yu et al. (2025) , leaf, text width=62.0em ] ] ] [ Training
@@
 (§3.5), fill=yellow!20 [ Alignment (§3.5), fill=red!15 [ VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al. (2024),  Zhi Lim et al. (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al. (2024) , leaf, text width=62.0em ] ] [ Robustness (§D.1), fill=red!15 [  Buettner and Kovashka (2024), MORE Cui et al. (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al. (2024), RA-BLIP Ding et al. (2024b), RagVL Chen et al. (2024e), RA-CM3 Yasunaga et al. (2023) , leaf, text width=62.0em ] ] ] ] ]
@@ 3.1 Retrieval Strategy
Efficient Search and Similarity Retrieval
Modern multimodal RAG systems encode diverse input modalities into a unified embedding space to enable direct cross-modal retrieval. Early CLIP-based Radford et al. (2021) methods often struggled to balance retrieval precision and computational cost. BLIP-inspired Li et al. (2022) approaches addressed some of these trade-offs by integrating cross-modal attention during training, yielding richer alignments between visual and textual features. To reconcile high accuracy with efficiency, contrastive retrieval frameworks such as MARVEL Zhou et al. (2024c) and Uni-IR Wei et al. (2024a) improved inter-modal discrimination through hard-negative mining and balanced sampling strategies Zhang et al. (2024i); Lan et al. (2025). Despite these representational gains, direct search over millions of embeddings demands fast similarity computation. Maximum inner product search (MIPS) variants offer sublinear lookup by approximating top-$k$ inner products Tiwari et al. (2024); Wang et al. (2024c); Zhao et al. (2023b). However, coarse quantization can degrade recall. To mitigate this, adaptive quantization methods Zhang et al. (2023a); Li et al. (2024a) dynamically allocate bits where the embedding distribution is dense, resulting in recall improvements over static schemes. Hybrid sparse–dense retrieval Nguyen et al. (2024); Zhang et al. (2024a) further complements dense embeddings with sparse lexical signals. Systems such as MuRAG Chen et al. (2022a) and RA-CM3 Yasunaga et al. (2023) employ approximate MIPS for efficient top-k candidate retrieval from large collections of image–text embeddings. Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al. (2022), for high-speed retrieval. Other efficient similarity computation methods include ScaNN (Scalable Nearest Neighbors) Guo et al. (2020), MAXSIM score Chan and Ng (2008); Cho et al. (2024), and approximate KNN methods Caffagni et al. (2024). Emerging approaches explore learned index structures Zhai et al. (2023); Basnet et al. (2024), which embed the search tree itself in neural parameters, aiming to adapt retrieval paths to the data distribution and reduce both latency and storage overhead.
Modality-Based Retrieval
Modality-aware retrieval techniques optimize efficiency by leveraging the unique characteristics of each modality. (i) Text-centric retrieval remains foundational in multimodal RAG systems, with both traditional methods like BM25 Robertson and Zaragoza (2009) and dense retrievers such as MiniLM Wang et al. (2020a) and BGE-M3 Chen et al. (2024b) dominating text-based evidence retrieval Chen et al. (2022b); Suri et al. (2024); Nan et al. (2024b). Novel approaches also address the need for fine-grained semantic matching and domain specificity: For instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al. (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al. (2024h) and CRAG Yan et al. (2024) enhance retrieval by ensuring accurate citation of text spans. (ii) Vision-centric retrieval leverages image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al. (2023). Systems such as EchoSight Yan and Xie (2024) and ImgRet Shohan et al. (2024) retrieve visually similar content by using reference images as queries. In addition, composed image retrieval methods Feng et al. (2023); Zhao et al. (2024); Jang et al. (2024); Saito et al. (2023) integrate multiple image features into unified query representations, enabling zero-shot image retrieval. (iii) Video-centric retrieval extends vision-based techniques by incorporating temporal dynamics and large video-language models. For instance, iRAG Arefeen et al. (2024) enables incremental retrieval for sequential video understanding, addressing the need for temporal coherence, while T-Mass Wang et al. (2024b) uses stochastic text embeddings to improve robustness in text-video alignment. Tackling long-context processing, Video-RAG Luo et al. (2024b) avoids reliance on proprietary models by using auxiliary OCR/ASR texts, whereas VideoRAG Ren et al. (2025) employs dual-channel architectures and graph-based knowledge grounding for extreme-length videos. To capture temporal reasoning, CTCH Shen et al. (2024) applies contrastive transformer hashing for long-term dependencies, which RTime Du et al. (2024) further refines by introducing reversed-video hard negatives for more robust causality benchmarking. Finally, OmAgent Zhang et al. (2024e) addresses the challenge of complex video understanding with a divide-and-conquer framework, while DRVideo Ma et al. (2024e) takes a complementary document-centric approach to enhance narrative preservation. (iv) Audio-centric retrieval aims to bypass traditional ASR pipelines while improving contextual alignment and real-time processing Xue et al. (2024a); Ghosh et al. (2024); Min et al. (2025). Pioneering frameworks like WavRAG Chen et al. (2025b) and SEAL Sun et al. (2025) introduce unified embedding architectures, directly mapping raw audio into a shared latent space to enable retrieval from hybrid knowledge bases. Audiobox TTA-RAG Yang et al. (2024a) conditions text-to-audio synthesis on retrieved acoustic samples, thereby enhancing zero-shot performance by enriching prompts with unlabeled audio context. For audio captioning, DRCap Li et al. (2025c) bridges the audio-text latent space of CLAP Wu et al. (2023) via text-only training for domain-adaptable descriptions without paired data. In parallel, P2PCAP Changin et al. (2024) improves retrieval precision by regenerating captions as dynamic queries. Further innovations address error correction and efficiency. LA-RAG Li et al. (2024b) utilizes fine-grained speech-to-speech retrieval and forced alignment to enhance ASR accuracy through LLM in-context learning. Meanwhile, hybrid systems, such as Xiao et al. (2025), integrate LLMs to correct errors in noisy environments using retrieved text/audio context.
@@
Recent research has moved beyond traditional unimodal retrieval, developing models that process entire documents by integrating textual, visual, and layout information. ColPali Faysse et al. (2024) pioneers end-to-end document image retrieval by embedding page patches with a vision-language backbone, bypassing OCR entirely. Models like ColQwen2 Wang et al. (2024d); Faysse et al. (2024) and M3DocVQA Cho et al. (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning. Newer frameworks refine efficiency and layout understanding: ViTLP Mao et al. (2024) and DocLLM Wang et al. (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al. (2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computational costs. Finally, mPLUG-DocOwl 1.5 Hu et al. (2024a) and 2 Hu et al. (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies, while SV-RAG Chen et al. (2025a) leverages MLLMs’ intrinsic retrieval capabilities via dual LoRA adapters: one for evidence page retrieval and the other for question answering.
Re-ranking and Selection Strategies
Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also prioritizing retrieved candidates. Re-ranking and selection strategies improve retrieval quality through optimized example selection, refined relevance scoring, and filtering mechanisms. (i) Optimized example selection techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection approaches Luo et al. (2024a); Yuan et al. (2023). Supervised methods like Su et al. (2024a) enhance multimodal inputs using probabilistic control keywords, whereas RULE Xia et al. (2024b) calibrates retrieved context via statistical methods like the Bonferroni correction Haynes (2013) to mitigate factuality risks. Clustering-based key-frame selection ensures diversity in video-based retrieval Dong et al. (2024b). Advanced (ii) scoring mechanisms are employed by several methods to improve retrieval relevance Mortaheb et al. (2025b, a); Zhi Lim et al. (2024). Multimodal similarity measures, including structural similarity index measure (SSIM) Wang et al. (2020b), normalized cross-correlation (NCC), and BERTScore Zhang et al. (2020), aid in re-ranking documents. Some frameworks combine similarity scores derived from various modalities for more robust re-ranking. For example, VR-RAG Khan et al. (2025) proposes a visual re-ranking framework that combines cross-modal text-image similarity with intra-modal visual similarity using DINOv2 Oquab et al. (2023), demonstrating significant improvements in open-vocabulary recognition tasks. Hierarchical post-processing integrates passage-level and answer confidence scores for improved ranking Zhang et al. (2024g); Yan and Xie (2024); Xu et al. (2024a). LDRE Yang et al. (2024b) employs semantic ensemble methods to adaptively weigh multiple caption features, while RAGTrans Cheng et al. (2024) and OMG-QA Nan et al. (2024b) incorporate traditional ranking functions like BM25 Robertson and Zaragoza (2009). (iii) Filtering methods ensure high-quality retrieval by eliminating irrelevant data. Hard negative mining, as used in GME Zhang et al. (2024i) and MM-Embed Lin et al. (2024a), mitigates modality bias through modality-aware sampling and synthesized negatives. Similarly, consensus-based filtering, seen in MuRAR Zhu et al. (2025) and ColPali Faysse et al. (2024), employs source attribution and multi-vector mapping to filter out low-similarity candidates. Dynamic modality filtering methods, such as RAFT Zhang et al. (2024h) and MAIN-RAG Chang et al. (2024), train retrievers to disregard confusing data, improving multimodal retrieval robustness.